# EfficientQAT: Efficient Quantization-Aware Training for Large Language Models

### v2

@@ 4.2 EfficientQAT for Instruction Tuning
Table 6: W2g64 Llama-2-7B performance with different trainable parameters in the block-wise training (w/o E2E-QP). “# Param.” indicates trainable parameters count in a block.
Param. # Param. Memory Avg. PPL Avg. Accuracy
@@
4.3 Ablation Analysis
Trainable parameters of Block-AP. Block-AP trains all parameters, including original weights and quantization parameters. Previous methods have introduced various training strategies to mitigate overfitting, such as trained rounding (Nagel et al., 2020; Cheng et al., 2023), clipping thresholds (Shao et al., 2023), and step sizes (Esser et al., 2019; Ding et al., 2023). We compare Block-AP with these methods by modifying only the trainable parameters of Block-AP. As shown in Table 7, Block-AP (training $s$, $z$, $\mathbf{W}$) performs best with an acceptable training cost. Additionally, the memory footprint of directly training $\mathbf{W}$ is even smaller than that of training the rounding operation, which requires an additional copy of rounding parameters.
@@ 4.3 Ablation Analysis
Figure 4: Llama-2-7B w2g64 quantization performance with different sample numbers for E2E-QP (w/ Block-AP).
@@
Llama-2 Block-AP E2E-QP
Time Memory Time Memory (4-/3-/2-bits) Total Time
7B 3.3h 8.5GB $\sim$1.5h 7.0/6.4/5.6GB 4.8h
13B 5.6h 10.3GB $\sim$2.9h 11.7/10.6/9.1GB 8.5h
70B 26.6h 29.9GB $\sim$14.3h 48.4/42.0/34.2GB 40.9h
@@
4.4 Efficiency of EfficientQAT
Training Efficiency Table 8 illustrates the required memory and time for training Lllama-2 models using EfficientQAT. The results indicate that the model completes training rapidly, taking 4.8 hours for the 7B model and 40.9 hours for the 70B model. we further compare the training time with other QAT methods, including BitDistiller, PB-LLM, and DN-LLM. As shown in Table 1(c), the training time of EfficientQAT is significantly lower than that of existing methods. For example, the tuning time of EfficientQAT is only 14% of AQLM and 50% of DB-LLM. Additionally, for quantizing a 70B model, the full process of EfficientQAT can be completed on a single A100-80GB GPU. However, other methods, except AQLM, require at least 4 A100-80GB GPUs to quantize a model of this size. Therefore, EfficientQAT is both a time-efficient and memory-efficient QAT method.
@@ Appendix C Speedup with BitBlas
In this study, we present a thorough comparison of our method against existing PTQ techniques, including GPTQ (Frantar et al., 2022), AWQ (Lin et al., 2023), OmniQ (Shao et al., 2023), AutoRound (Cheng et al., 2023), QuIP# (Tseng et al., 2024), and AQLM (Egiazarian et al., 2024). We also compare with existing QAT methods, including LLM-QAT (Liu et al., 2023e), BitDistiller (Du et al., 2024), PB-LLM (Shang et al., 2023) and DB-LLM (Chen et al., 2024). Additionally, we also evaluate quantized parameter-efficient fine-tuning methods such as PEQA (Kim et al., 2023a), QLoRA (Dettmers et al., 2023a), QA-LoRA (Xu et al., 2023b), and IR-QLoRA (Qin et al., 2024). The results we discuss originate from their respective official publications, and other scholarly articles, or are derived from our reproduction. We meticulously document the source of the results for each method as follows:
@@
  • ApiQ: These results are cited from IR-ApiQ (Liao & Monz, 2024) paper.
@@ Appendix F Additional Ablation Analysis
Appendix H Full Results
In Table 1, we present the average accuracy for five zero-shot tasks. This section offers a detailed breakdown of the task-specific accuracy numbers. Specifically, Tables 15, 16, and 17 detail the performance of 4-bit, 3-bit, and 2-bit quantization, respectively.
Table 15: 4-bit Llama 2 & 3 zero-shot accuracy by lm_eval v0.4.2 ( acc is reported, not acc_norm )
2-7B - - 16 69.22 57.16 43.52 76.26 78.07 64.85
RTN 4 128 68.35 56.91 43.52 76.26 77.58 64.52
GPTQ 4 128 69.06 56.36 42.15 75.63 78.02 64.24
AWQ 4 128 68.98 56.40 43.86 76.14 77.31 64.54
OmniQ 4 128 68.98 56.59 43.34 75.76 77.91 64.52
AutoRound 4 128 68.67 56.79 42.58 75.76 78.13 64.39
QuIP# 4 - 69.22 56.65 43.00 75.51 78.02 64.48
EfficientQAT 4 128 69.22 57.00 42.32 75.13 77.69 64.27
2-13B - 16 - 72.22 60.07 48.29 79.42 79.05 67.81
RTN 4 128 72.14 59.77 47.95 79.00 78.62 67.50
GPTQ 4 128 72.14 59.76 47.53 78.58 78.35 67.27
AWQ 4 128 73.32 59.80 46.50 79.38 79.05 67.61
OmniQ 4 128 72.06 59.53 47.18 78.37 78.35 67.10
AutoRound 4 128 71.67 59.87 47.01 79.25 79.00 67.36
QuIP# 4 - 72.69 59.49 46.59 78.62 79.00 67.28
EfficientQAT 4 128 71.98 59.87 47.53 79.34 78.89 67.52
2-70B - 16 - 77.98 64.77 54.44 82.70 82.15 72.41
RTN 4 128 78.14 63.93 54.78 82.79 81.66 72.26
GPTQ 4 128 78.22 64.45 54.61 82.79 81.88 72.39
AWQ 4 128 77.58 64.48 55.12 82.70 82.32 72.44
OmniQ 4 128 77.51 64.52 55.12 82.91 81.88 72.39
AutoRound 4 128 78.30 64.60 54.52 82.87 82.05 72.47
QuIP# 4 - 77.82 64.51 54.44 82.37 81.72 72.17
EfficientQAT 4 128 78.45 64.57 55.12 83.00 81.94 72.62
3-8B - - 16 72.61 60.17 50.43 80.09 79.60 68.58
RTN 4 128 73.16 59.04 48.38 79.25 79.11 67.79
GPTQ 4 128 73.72 59.17 47.78 79.38 78.94 67.80
AWQ 4 128 73.01 59.43 50.00 79.55 79.22 68.24
EfficientQAT 4 128 72.53 59.43 50.94 79.84 79.43 68.43
3-70B - 16 80.51 66.36 60.41 86.99 82.37 75.33
RTN 4 128 78.77 65.83 57.34 85.69 82.26 73.98
GPTQ 4 128 80.51 66.12 59.04 85.77 82.26 74.74
AWQ 4 128 80.35 65.82 59.13 86.41 82.15 74.77
EfficientQAT 4 128 79.24 66.27 59.13 85.86 82.37 74.57

### v3

@@ 4.2 EfficientQAT for Instruction Tuning
Table 2: Llama 2 & 3 Wikitext2 and C4 perplexity ($\downarrow$), context length 2048. "-" indicates the result is unreachable in the public papers.
Wikitext 2 C4
Method Bits Group 2-7 2-13 2-70 3-8 3-70 2-7 2-13 2-70 3-8 3-70
FP16 16 - 5.47 4.88 3.32 6.14 2.85 6.97 6.47 5.52 8.88 6.73
GPTQ 3 128 6.29 5.42 3.85 9.58 5.25 7.89 7.00 5.85 11.66 8.64
AWQ 3 128 6.24 5.32 3.74 8.16 4.69 7.84 6.94 5.81 11.49 7.91
OmniQ 3 128 6.03 5.28 3.78 8.27 4.99 7.75 6.98 5.85 11.66 7.97
BitDistiller 3 128 5.97 - - - - - - - - -
EfficientQAT 3 128 5.81 5.12 3.61 7.09 4.21 7.34 6.73 5.71 10.06 7.46
OmniQ 2 128 11.06 8.26 6.55 18.50 16.79 15.02 11.05 8.52 22.46 15.06
ApiQ 2 128 8.25 6.71 - - - 12.04 9.13 - - -
BitDistiller 2 128 8.08 - - - - - - - - -
EfficientQAT 2 128 7.19 6.08 4.61 9.80 6.38 8.79 7.75 6.48 13.22 9.53
AQLM 2 2x8 7.24 6.06 4.49 - - 8.96 7.80 6.36 - -
QuIP# 2 - 6.66 5.74 4.16 - - 8.35 7.45 6.12 - -
ApiQ 2 64 7.59 6.44 - - - 10.56 8.92 - - -
CBQ 2 64 8.01 - - - - 11.30 - - - -
DB-LLM 2 64 7.23 6.19 4.64 13.60 - 9.62 8.38 6.77 19.20 -
EfficientQAT 2 64 6.86 5.96 4.52 9.41 6.07 8.50 7.59 6.38 12.77 9.23
@@
4.3 Ablation Analysis
Trainable parameters of Block-AP. Block-AP trains all parameters, including original weights and quantization parameters. Previous methods have introduced various training strategies to mitigate overfitting, such as trained rounding (Nagel et al., 2020; Cheng et al., 2023), clipping thresholds (Shao et al., 2023), and step sizes (Esser et al., 2019; Ding et al., 2023). We compare Block-AP with these methods by modifying only the trainable parameters of Block-AP. As shown in Table 5, Block-AP (training $s$, $z$, $\mathbf{W}$) performs best with an acceptable training cost. Additionally, the memory footprint of directly training $\mathbf{W}$ is even smaller than that of training the rounding operation, which requires an additional copy of rounding parameters. Additionally, BitNet Ma et al. (2024) demonstrates that optimizing only the weights, without considering quantization parameters, can still achieve strong performance. However, Table 5 shows that training only the weights results in a perplexity of 14.32, which is significantly higher than the 8.53 achieved by Block-AP. This difference arises because our quantization approach starts from a pre-trained model and directly optimizes the scaling factors (s) and zero points (z) to minimize quantization errors, making minimal changes to the weights and thus preserving the model’s learned knowledge. In contrast, training only the weights adjusts the scaling factors indirectly, requiring larger weight updates that can disrupt this knowledge. BitNet Ma et al. (2024), which is trained from scratch, does not face this issue.
@@ 4.3 Ablation Analysis
Table 5: W2g64 Llama-2-7B performance with different trainable parameters in the block-wise training (w/o E2E-QP). “#” indicates trainable parameters count in a block.
Param. # Memory Avg. PPL Avg. Acc.
@@
Llama-2 Block-AP E2E-QP
Time Memory Time Memory (4-/3-/2-bits) Total Time
7B 3.3h 8.5GB $\sim$1.5h 7.0/6.4/5.6GB 4.8h
13B 5.6h 10.3GB $\sim$2.9h 11.7/10.6/9.1GB 8.5h
70B 26.6h 29.9GB $\sim$14.3h 48.4/42.0/34.2GB 40.9h
@@
4.4 Efficiency of EfficientQAT
Training Efficiency Table 7 illustrates the required memory and time for training Lllama-2 models using EfficientQAT. The results indicate that the model completes training rapidly, taking 4.8 hours for the 7B model and 40.9 hours for the 70B model. we further compare the training time with other QAT methods, including BitDistiller, and DB-LLM. As shown in Table 9, the training time of EfficientQAT is significantly lower than that of existing methods. For example, the tuning time of EfficientQAT is only 50% of DB-LLM. Additionally, for quantizing a 70B model, the full process of EfficientQAT can be completed on a single A100-80GB GPU. However, other methods require at least 4 A100-80GB GPUs to quantize a model of this size. Therefore, EfficientQAT is both a time-efficient and memory-efficient QAT method.
Table 8: Llama-2-7B w2g64 quantization performance with different sample numbers for E2E-QP (w/ Block-AP).
@@ 5 Conclusion
6 Limitation
EfficientQAT achieves impressive results in low-bit quantization scenarios, but there remains a performance gap compared to full-precision (FP16) models, particularly in 2-bit settings. Reducing this gap without sacrificing efficiency remains a challenge. Additionally, the method depends on the availability of high-quality and diverse datasets, requiring 4096 samples for effective training in both the Block-AP and E2E-QP phases. The performance of the quantized models can vary significantly based on the size and distribution of the training data. This reliance may limit its effectiveness in data-scarce or domain-specific applications.
Acknowledgement
This paper is partially supported by the National Key R&D Program of China No.2022ZD0161000.
@@ Appendix C Speedup with BitBlas
In this study, we present a thorough comparison of our method against existing PTQ techniques, including GPTQ (Frantar et al., 2022), AWQ (Lin et al., 2023), OmniQ (Shao et al., 2023), AutoRound (Cheng et al., 2023), QuIP# (Tseng et al., 2024), and AQLM (Egiazarian et al., 2024). We also compare with existing QAT methods, including LLM-QAT (Liu et al., 2023e), BitDistiller (Du et al., 2024), PB-LLM (Shang et al., 2023) and DB-LLM (Chen et al., 2024a). Additionally, we also evaluate quantized parameter-efficient fine-tuning methods such as PEQA (Kim et al., 2023a), QLoRA (Dettmers et al., 2023a), QA-LoRA (Xu et al., 2023b), and IR-QLoRA (Qin et al., 2024). The results we discuss originate from their respective official publications, and other scholarly articles, or are derived from our reproduction. We meticulously document the source of the results for each method as follows:
@@
  • ApiQ: These results are cited from IR-ApiQ (Liao and Monz, 2024) paper.
@@ Appendix F Additional Ablation Analysis
Appendix H Comparisons with the Same Number of Data Samples
The main experiments use 4096 samples for the proposed method. However, some PTQ methods, such as OmniQuant Shao et al. (2023) and GPTQ Frantar et al. (2022), use only 128 samples for quantization. To ensure a fair comparison, we also evaluate EfficientQAT against OmniQuant and GPTQ using the same number of data samples. As shown in Table 15, the performance of OmniQuant Shao et al. (2023) and GPTQ Frantar et al. (2022) stabilizes at 128 samples and does not improve with additional data, while EfficientQAT continues to benefit from more samples. Even with only 128 samples, EfficientQAT significantly outperforms OmniQuant (8.02 PPL vs. 15.02 PPL). Furthermore, Table 1 shows that EfficientQAT surpasses DB-LLM, which uses 20k samples, despite EfficientQAT using only 4096 samples. These results confirm the consistent superiority of EfficientQAT over other uniform quantization methods, highlighting its effectiveness.
Method Precision 64 128 256 512
GPTQ W3g128 7.91 7.89 7.90 7.89
OmniQuant W3g128 7.70 7.75 7.73 7.74
EfficientQAT W3g128 7.40 7.37 7.36 7.35
OmniQuant W2g128 15.23 15.02 14.95 14.93
EfficientQAT W2g128 9.01 8.95 8.85 8.83
Table 15: C4 perplexity of Llama-2-7B with different training samples.
Appendix I Full Results
In Table 1, we present the average accuracy for five zero-shot tasks. This section offers a detailed breakdown of the task-specific accuracy numbers. Specifically, 16 and 17 detail the performance of 3-bit and 2-bit quantization, respectively.